Today, Belle's age is four less than twice the age of her sister, Cecille. in five years, twice Belle's age will be three times Cecille age. How old are they today?
Answer by lwsshak3(11628)

Today, Belle's age is four less than twice the age of her sister, Cecille. in five years, twice Belle's age will be three times Cecille age. How old are they today?
***
Today:
let x=Cecille's age
2x-4=Belle's age
..
In 5 yrs:
x+5=Cecille's age
2x-4+5=2x+1=Belle's age
..
2(2x+1)=3(x+5)
4x+2=3x+15
x=13
2x-4=22
..
Today:
Cecille's age=13
Belle's age=22